The tires of a car make 95 revolutions as the car reduces its speed uniformly from 95.0 km/hkm/h to 55.0 km/hkm/h. The tires have a diameter of 0.88 m

Angular acceleration of the tires: -2 rad/s2

A. If the car continues to decelerate at this rate, how much more time is required for it to stop?

Express your answer to two significant figures and include the appropriate units.

B. If the car continues to decelerate at this rate, how far does it go? Find the total distance.

Express your answer to three significant figures and include the appropriate units.
Answer #1

u=95 km/h = 26.4 m/s U = 55kmur = 15.3 m/s &=-2 nads a=- 2 x 0.44 18 .: Accelenation =-0.88m 1st A Y = 0, U= 15.30m/s, a = -0